# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v4

@@ 8 Container Density
Now that we have established interior regularity and interior degree doubling, we can start to think about what those things mean. The Weak Exteriors Decreasing Lemma implies that containers will decrease in size. We need to see how that decrease in size will impact the degrees of the nodes.
@@
Regular Exterior Degrees Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node. Let a minimum degree node of $G$ be $v_{0}$, and suppose every node in $G$ has the Decreasing Sequence Property (DSP). Then the regular interior degrees will result in regular exterior degrees in $G$.
@@
Let $x_{i}$ and $x_{j}$ be two nodes at the same distance $i$ from the minimum degree node $v_{0}$ in the oriented graph $G$, both contained in the same container $C_{i}$.
@@
Since the size of the next container $C_{i+1}$ is determined by the exterior degrees of nodes in $C_{i}$, this irregularity in the exterior degrees would force the size of $C_{i+1}$ to differ, contradicting the assumption that the containers are of regular size.
By Lemmas 6.1 (Regular Interior Degrees) and 7.2 (Interior Degree Double), we know that if the interior degrees within a container are regular, the exterior degrees must also follow this regularity. Specifically, if there is an irregularity, any parent node of $x_{i}$ and $x_{j}$ will have its exterior degree double, which would in turn double their total degree.
This doubling of the parents’ total degree would violate the Decreasing Sequence Property (DSP).
Therefore, the assumption that $|ext(x_{i},y)|$ and $|ext(x_{j},y)|$ differ is false. Hence, the exterior degrees in $G$ must be regular if the interior degrees are regular. ∎
We proved Lemma 6.1 (Regular Interior Degrees) on our way to proving that back arcs did not exist. LABEL:refintdeg is an important property of the graph though that we are going to use to get other properties of the graph, like 8.1.
@@
Regular Graph All Nodes Have Equal Our-Degree Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node, with $v_{0}$ as a minimum degree node. Further, let all nodes satisfy the Decreasing Sequence Property (DSP). Then for any node $v$ in $G$, the degree of $v$ is equal to $d$, where $d_{v_{0}}=d^{+}(v_{0})$.
Proof.
We will prove this lemma by induction on the number of nodes in the graph $G$.
For the base case, consider the node $v_{0}$, which, by definition, has the minimum degree $d$, so $d^{+}(v_{0})=d_{v_{0}}$.
Assume that for some node $v_{k}$ in $G$, the degree of $v_{k}$ satisfies $d^{+}(v_{k})=d_{v_{0}}$.
Now consider an arbitrary node $v_{k+1}$ in $G$. Suppose, for contradiction, that $d^{+}(v_{k+1})>d_{v_{0}}$.
Consider a neighbor $v_{n}$ of $v_{k+1}$. By the inductive hypothesis, the degree of $v_{n}$ is exactly $d_{v_{0}}$. Since $v_{k+1}$ has more than $d_{v_{0}}$ neighbors, the neighbors of $v_{k+1}$ are second neighbors of $v_{n}$ in the graph $G^{2}$. Therefore, $v_{n}$ would have more than $d_{v_{0}}$ second neighbors in $G^{2}$.
This implies that $v_{n}$’s degree has effectively doubled, which contradicts the DSP, which requires that $v_{n}$ must have degree $d_{v_{0}}$. Hence, the assumption that $d^{+}(v_{k+1})>d_{v_{0}}$ leads to a contradiction.
Thus, we must conclude that:
$d^{+}(v_{k+1})=d_{v_{0}}.$
By the principle of mathematical induction, it follows that for every node $v$ in $G$, the degree of $v$ is equal to $d_{v_{0}}$. ∎
Lemma 8.3.
Strong Decreasing Exteriors Lemma Let $G$ be an oriented graph represented by a GLOVER In an attempt to prevent a degree doubling node, with no back arcs, and let $v_{0}$ be the chosen minimum degree node in $G$. For any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of the exterior neighbors must be bounded by
@@
for some $z\in N^{+}(x_{i})$ in order to keep any node’s degree from doubling in the oriented graph.
@@
Base Case ($i=1$): Consider a first neighbor $x_{1}\in C_{1}$ (the nodes at distance 1 from $v_{0}$). The size of the exterior set
@@
is bounded by $d^{+}(v_{0})-1$, since each first neighbor $x_{1}$ can connect to at most $d^{+}(v_{0})-1$ nodes at distance 2, as there are no back arcs to $v_{0}$.
Inductive Hypothesis: Assume that for any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of its exterior neighbors is bounded by $d^{+}(v_{0})-(i-1)$. Specifically, we assume:
@@
Lemma (Strong Decreasing Exteriors Lemma) uses Lemma 5.6 (Weak Decreasing Exteriors Lemma), to give us a way of calculating the maximum size of a container based only in its distance from the minimum degree node and the minimum degree node’s degree.
The immediate takeaway from this lemma is the decreasing sequence of exteriors that we were attempting to prove. It is for this sequence that the term ”decreasing sequence property” is named.
@@
Interior Degrees Lower Bound Let $G$ be an oriented graph represented by a GLOVER, In an attempt to prevent a degree doubling node, and let $v_{0}$ be a minimum degree node in $G$, where every node in $G$ has the Decreasing Sequence Property (DSP). Let $x$ be a node at distance $i$ from $v_{0}$, and $v_{i-1}$ be a node at distance $i-1$ from $v_{0}$. Then the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$, provided $(v_{0},x)\in G$.
@@
Now, since $x$ has at least the same out-degree as $v_{0}$ (by the assumption of DSP), $x$ must map to at least $i$ other nodes. Given that back arcs are not allowed (as shown previously), $x$ cannot map to nodes at earlier distances without violating the structure of the graph.
Therefore, the only remaining possibility is that $x$ maps to interior nodes. Since $x$’s degree must account for at least $i$ mappings, and these must be to interior nodes, we conclude that $x$ must map to at least $i$ interior nodes.
Thus, $|int(v_{i-1},x)|\geq i$, as required. ∎
Having already placed an upper bound on the exterior sizes and containers, it is only natural that the there is an opposite bound for the interiors. Once again the minimum degree node is the driving force behind this. The degree of these interior nodes will also be impacted by their distance to the minimum degree node.
@@ 9 Transitive Triangles and Seymour Diamonds
Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, $u$ form a transitive triangle.
Here, we have $x$ pointing to $y$ and $u$; and $y$ pointing to $u$. The concept of a transitive triangle is an important one because it involves the risk of double counting. The node $u$ is both a first neighbor of $x$ and a second neighbor of $x$, so if we do not properly deal with transitive triangles we might improperly count out second neighbors and thus not have a correct algorithm.
@@
Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
The Seymour Diamond is a graph theory concept in graph there where a node $x$ points to two out-neighbors $y$ and $w$, and these two out-neighbors both share the same out-neighbor, $u$. This is basically a four vertex oriented graph. What this situation presents is redundant second neighbors. There is a path $x\rightarrow u_{1}\rightarrow u_{2}$ in this diamond, and there is a path $x\rightarrow u_{2}\rightarrow u_{1}$ in this diamond. As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and more importantly the Seymour’s Second Neighborhood Conjecture.
The notion of interior neighbors is another way of describing what [1] referred to as transitive triangles. They showed that these triangles are always present in a graph that doesn’t satisfy Seymour’s conjecture, i.e. one where every node has the decreasing sequence property. We provided a similar result with Lemma 5.4 (Interior Degree Greater than 1). When a minimum degree node $v_{0}$ has the decreasing sequence property, all of its neighbors have interior degree of at least 1. That means that every node $x\in N^{+}(v_{0})$ will relate to some other node $y\in N^{+}(v_{0})$ and these three nodes, $(v_{0},x,y)$ will form a transitive triangle. We can extend that further for any two constructive containers.
@@
Transitive Triangles from Containers In a graph $G$ represented by a GLOVER, for any two consecutive containers $C_{i}$ and $C_{i+1}$ in a graph satisfying the Decreasing Sequence Property (DSP), there exist two parent nodes $u\in C_{i}$ and $v\in C_{i}$, and a shared child node $w\in C_{i+1}$, such that the nodes $u$, $v$, and $w$ form a transitive triangle. That is, there exist directed edges from $u$ to $w$, $v$ to $w$, and from $u$ to $v$, or from $v$ to $u$.
@@
In a graph $G$ represented by a GLOVER, for any three consecutive containers $C_{i}$, $C_{i+1}$, and $C_{i+2}$, let:
@@
The following arcs hold:
1. There are forward arcs $x\to y$ and $y\to u$, 2. There are forward arcs $x\to w$ and $w\to u$, 3. No back arcs exist between containers.
Similar to Lemma 9.1, Lemma 9.2 (Seymour Diamonds from Containers) lets us know that we have been interacting with Seymour Diamonds all along without knowing it. We didn’t know it because they they did not disrupt our algorithm and they are not a hindrance to our algorithm.
@@
Seymour Diamonds and Decreasing Path Algorithm Let $G$ be an oriented graph represented by the GLOVER, in an attempt to prevent a degree during node, where we are applying the DSP path algorithm. Suppose $G$ contains a subgraph $D$ that forms a Seymour diamond. Then, the presence of the Seymour diamond $D$ does not disrupt the DSP path algorithm.
@@
  • The DSP path algorithm can utilize one of these paths while disregarding the other, as both connect $x$ and $u$, thus maintaining degree constraints.
  • The presence of redundant paths does not diminish the outgoing neighbors of any node, ensuring that the minimum degree condition holds throughout the process.
  • Consequently, these additional connections offer alternative routes without introducing cycles or otherwise disrupting the algorithm’s operation.
None of these case are a deterrent to the execution of the DSP Path algorithm. ∎
@@
Interaction of Minimum Degree Nodes and Transitive Triangles Let $G$ be an oriented graph represented by a GLOVER, containing a minimum degree node $v_{0}$ with degree $d$. Consider a transitive triangle formed by nodes $z$, $y$, and another vertex $x$, where $z$ and $y$ are part of this triangle. The following cases are possible:
  • If $d^{+}(y)=d$ and $d^{+}g(z)=d$
  • If $d^{+}(y)=d$ and $d^{+}(z)>d$
  • If $d^{+}(y)>d$
  • If $d^{+}(y)<d$
@@
  • Case 1: If $d^{+}(y)=d$ and $d^{+}(z)=d$: Both $y$ and $z$ have exactly $d$ outgoing connections. For the triangle to be transitive, they must not share all neighbors, meaning there exists at least one neighbor of $y$ that is not a neighbor of $z$. Thus, $N^{+}(y)\nsubseteq N^{++}(z)$.
  • Case 2: If $d^{+}(y)=d$ and $d^{+}(z)>d$: $z$ covers some neighbors of $y$, but since $y$ must map to at least one neighbor unique from those of $z$, it follows that $y$ has a neighbor that $z$ does not connect to.
  • Case 3: If $d^{+}(y)>d$: $y$ has more outgoing connections than required. It can utilize these additional connections to map to unique neighbors without impacting the minimum degree condition, ensuring $y$ retains neighbors not covered by $z$.
  • Case 4: If $d^{+}(y)<d$: This scenario contradicts the minimum degree requirement since each node must have at least $d$ outgoing connections. Therefore, $y$ cannot be part of the triangle involving $v_{0}$.
Now that we see that transitive triangles and Seymour diamonds are not a problem for our algorithm, we can continue our investigation of oriented graphs. Next we will blend everything together into our main results.
@@ 10.1 Introduction
Here is where we sum up all the lemmas we have derived from our analysis on oriented graphs. We will state the final algorithm and some final results.
So far, this battle has been a battle between a catalyst - the chosen minimum degree node, and the bottleneck - the Decreasing Sequence Property (DSP). We have proved some important things - Back arcs cannot exist; interior degrees must double inside each container, the exterior degrees decrease in size; interior degrees proportionally increase in size.
There is a limit though, to how large the interior degree can grow. It is bound by $\frac{n-1}{2}$ by the limit of average degree inside a graph.
@@
Let $G$ be an oriented graph represented by a GLOVER, in an attempt to prevent a degree doubling node, with all nodes satisfying the Decreasing Sequence Property (DSP), where $v_{0}$ is a minimum degree node in $G$, and all nodes are placed into containers. If there are an back arcs in $G$, then a node’s degree will double.

### v6

@@ 8 Container Density
Now that we have established interior regularity and interior degree doubling, we can start to think about what those things mean. The Weak Decreasing Exteriors Lemma implies that containers will decrease in size. We need to see how that decrease in size will impact the degrees of the nodes.
@@
Regular Exterior Degrees
Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree-doubling node. Let a minimum degree node of $G$ be $v_{0}$, and suppose every node in $G$ has the Decreasing Sequence Property (DSP). Then the nodes in a container in $G$ will result in regular exterior degrees.
@@
Let $x_{i}$ and $x_{j}$ be two nodes in the same container $C_{i}$.
@@
By Lemmas 7.2 (Interior Degree Doubling) and 6.5 (No Back Arcs), we can imply Lemma 5.6 (Weak Decreasing Exteriors Lemma) here. This says that with each distance from the minimum degree node, our exterior size is decreased by 1.
This stopped at the container $C_{i}$. That means there will be two neighboring containers of equal size, $C_{i-1}$ and $C_{i}$. So, a node in $C-{i-2}$ can have its exterior degree double. By 7.2 and 7.1 (Indistinguishably of Regular Nodes), we know that this node has its interior degree double as well.
So, this assumption led to a contradiction.
This is an important property because it ties into the Decreasing Sequence Property (DSP). We were able to use the Weak Decreasing Exteriors Lemma in this proof because the two go hand in hand.
Anything that leads to a hiccup in the decrease in exteriors would cause a node’s degree to double. The DSP is a bottleneck in this way because it is really limiting what can take place in this graph.
@@
Strong Decreasing Exteriors Lemma
Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node in $G$. We are attempting to prevent a degree-doubling node, so for any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of the exterior neighbors must be bounded by
@@
for some $z\in N^{+}(x_{i})$.
@@
Base Case ($i=1$):
Consider a first neighbor $x_{1}\in C_{1}$ (the nodes at distance 1 from $v_{0}$). The size of the exterior set
@@
is bounded by $d^{+}(v_{0})-1$. Each first neighbor $x_{1}$ can connect to at most $d^{+}(v_{0})-1$ nodes at distance 2. We have already proved that there are no back arcs to $v_{0}$.
Inductive Hypothesis: Assume that for any node $x_{i}$ in container $C_{i}$, the size of its exterior neighbors is bounded by $d^{+}(v_{0})-(i-1)$. Specifically, we assume:
@@
Lemma 8.2 (Strong Decreasing Exteriors Lemma) uses Lemma 5.6 (Weak Decreasing Exteriors Lemma) to give us a way of calculating the maximum size of a container based only on its distance from the minimum degree node and the minimum degree node’s degree.
The immediate takeaway from this lemma is the decreasing sequence of exteriors that we were attempting to prove. It is for this sequence that the term ”decreasing sequence property” was originally named.
@@
Interior Degrees Lower Bound
Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node in $G$. Suppose every node in $G$ has the Decreasing Sequence Property (DSP). Let $x$ be a node at distance $i$ from $v_{0}$, and $v_{i-1}$ be a node at distance $i-1$ from $v_{0}$. Then the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$, provided $(v_{0},x)\in G$.
@@
Now, since $x$ has at least the same out-degree as $v_{0}$, $x$ must map to at least $i$ other nodes. Given that back arcs are not asked, $x$ cannot map to nodes at earlier distances.
Therefore, the only remaining possibility is that $x$ maps to interior nodes. Since $x$’s degree must account for at least $i$ mappings. These must be to interior nodes. So, we conclude that $x$ must map to at least $i$ interior nodes.
Thus, $|int(v_{i-1},x)|\geq i$, as required.
Having already placed an upper bound on the exterior sizes and containers, it is only natural that there is an opposite bound for the interiors. Once again, the minimum degree node is the driving force behind this. The degree of these interior nodes will also be impacted by their distance to the minimum degree node.
Lemma 8.4.
Regular Graph
All Nodes Have Equal Out-Degree
Let $G$ be an oriented graph represented by a GLOVER. Suppose $v_{0}$ is a minimum degree node of $G$. Further, let all nodes satisfy the Decreasing Sequence Property (DSP). We are attempting to prevent a degree-doubling node. Then for any node $v$ in $G$, the degree of $v$ is equal to $d$, where $\delta=d^{+}(v_{0})$.
Proof.
With the proofs on the lack of back arcs and the regularity of interior arcs and the regularity of exterior arcs, we arrive at the following equation for each container:
$\delta=|int(v_{p},u_{i})|+|ext(v_{p},u_{i})|$
where $u_{i}$ is in container $C_{i}$, and $v_{p}$ is a parent of $u_{i}$ in container $C_{i-1}$.
For a node $u_{i}$ in container $C_{i}$, with degree $\delta+1$, there would be the possible places for this degree to be located—back arcs, interior arcs, or exterior arcs. We have already shown that back arcs cannot exist. If this extra degree went to the exterior arcs, then it would cause the degree of a parent node to double.
The only remaining possibility is for this extra degree to go to the interior arcs of $u_{i}$. While this is not a problem, it will cause the interior degrees to increase at a faster rate than we just showed in Lemma 8.3 (Interior Degree Lower Bound), as this would increase the number of interior arcs.
This more rapid reduction in the interior degree leads to container density faster. To avoid this, regularity is key.
Thus, every node in every container’s degree is set by the minimum degree node. This leads to the regularity of the graph.
@@ 9 Transitive Triangles and Seymour Diamonds
Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, and $u$ form a transitive triangle.
Here, we have $x$ pointing to $y$ and $u$ and $y$ pointing to $u$. The concept of a transitive triangle is an important one because it involves the risk of double counting. The node $u$ is both a first neighbor of $x$ and a second neighbor of $x$, so if we do not properly deal with transitive triangles, we might improperly count out second neighbors and thus not have a correct algorithm.
@@
Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, and $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
The Seymour Diamond is a graph theory concept in graph theory where a node $x$ points to two out-neighbors $y$ and $w$, and these two out-neighbors both share the same out-neighbor, $u$. This is a four-vertex-oriented graph. What this situation presents is redundant second neighbors. There is a path $x\rightarrow u_{1}\rightarrow u_{2}$ in this diamond, and there is a path $x\rightarrow u_{2}\rightarrow u_{1}$ in this diamond. As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and, more importantly, the Seymour’s Second Neighborhood Conjecture.
The notion of interior neighbors is another way of describing what [1] referred to as transitive triangles. They showed that these triangles are always present in a graph that doesn’t satisfy Seymour’s conjecture, i.e., one where every node has the decreasing sequence property. We provided a similar result with Lemma 5.4 (Interior Degree at least 1). When a minimum degree node $v_{0}$ has the Decreasing Sequence Property, all of its neighbors have an interior degree of at least 1. That means that every node $x\in N^{+}(v_{0})$ will relate to some other node $y\in N^{+}(v_{0})$ and these three nodes, $(v_{0},x,y)$, will form a transitive triangle. We can extend that further for any two constructive containers.
@@
Transitive Triangles from Containers
Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP). For any two consecutive containers $C_{i}$ and $C_{i+1}$, there exist two parent nodes $u\in C_{i}$ and $v\in C_{i}$, and a shared child node $w\in C_{i+1}$, such that the nodes $u$, $v$, and $w$ form a transitive triangle.
@@
Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP). for any three consecutive containers $C_{i}$, $C_{i+1}$, and $C_{i+2}$, let:
@@
The following arcs exist:
1. There are forward arcs $x\to y$ and $y\to u$,
2. There are forward arcs $x\to w$ and $w\to u$,
3. No back arcs exist between containers.
Similar to Lemma 9.1, Lemma 9.2 (Seymour Diamonds from Containers) lets us know that we have been interacting with Seymour Diamonds all along without knowing it. We didn’t know it because they did not disrupt our algorithm.
@@ 2. There are forward arcs $x\to w$ and $w\to u$,
Seymour Diamonds and Decreasing Path Algorithm
Suppose $G$ is an oriented graph represented by the GLOVER, where all nodes have the Decreasing Sequence Property (DSP). We are attempting to prevent a degree-doubling node, where we are applying the DSP algorithm. Suppose $G$ contains a subgraph $D$ that forms a Seymour diamond. Then, the presence of the Seymour diamond $D$ does not disrupt the DSP algorithm.
@@
The DSP algorithm can utilize one of these paths while disregarding the other. Both paths connect $x$ and $u$, thus maintaining degree constraints.
The presence of redundant paths does not diminish the outgoing neighbors of any node. This ensures that the minimum degree condition holds throughout the process.
Consequently, these additional connections offer alternative routes without disrupting the algorithm’s operation.
None of these cases are a deterrent to the execution of the DSP algorithm.
@@
Interaction of Minimum Degree Nodes and Transitive Triangles
Suppose $G$ is an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node with degree $d$. Consider a transitive triangle formed by nodes $z$, $y$, and $x$, where $d^{+}(y)=1$ and $d^{+}(z)=0$.
Then the following cases are possible:
  • $d^{+}(y)=d$ and $d^{+}g(z)=d$
  • $d^{+}(y)=d$ and $d^{+}(z)>d$
  • $d^{+}(y)>d$
  • $d^{+}(y)<d$
@@
  • Case 1: $d^{+}(y)=d$ and $d^{+}(z)=d$:
    Both $y$ and $z$ have exactly $d$ outgoing connections. For the triangle to be transitive, they must not share all neighbors. This means that there exists at least one neighbor of $y$ that is not a neighbor of $z$. Thus, $N^{+}(y)\nsubseteq N^{++}(z)$.
  • Case 2: $d^{+}(y)=d$ and $d^{+}(z)>d$:
    Then $z$ may cover some neighbors of $N^{+}(y)$. Since $y$ must map to at least one neighbor unique from those of $z$, it follows that $y$ has a neighbor that $z$ does not connect to.
  • Case 3: If $d^{+}(y)>d$:
    $y$ has more outgoing connections than required. It can utilize these additional connections to map to unique neighbors without impacting the minimum degree condition, ensuring $y$ retains neighbors not covered by $z$.
  • Case 4: If $d^{+}(y)<d$:
    This scenario contradicts the minimum degree requirement since each node must have at least $d$ outgoing connections. Therefore, $y$ cannot be part of the triangle involving $v_{0}$.
Now that we see that transitive triangles and Seymour diamonds are not a problem for our algorithm, we can continue our investigation of oriented graphs. Next, we will blend everything together into our main results.
@@ 10.1 Introduction
Here, we will sum up all the lemmas we have derived from our analysis of oriented graphs. We will state the final algorithm and some final results.
So far, this battle has been a battle between a catalyst, the chosen minimum degree node and the bottleneck, the Decreasing Sequence Property (DSP). We have proved some important things: Back arcs cannot exist; interior degrees must double inside each container, and the exterior degrees decrease in size; interior degrees proportionally increase in size.
There is a limit, though, to how large the interior degree can grow. It is bound by $\frac{n-1}{2}$ by the limit of average degree inside a graph.
@@
Let $G$ be an oriented graph represented by a GLOVER, with all nodes satisfying the Decreasing Sequence Property (DSP). Let $v_{0}$ be a minimum degree node in $G$. We are attempting to prevent a degree-doubling node. If there are back arcs in $G$, then a node’s degree will double.